(x^3+y^3)=(3xy^2)dy/dx


1
Expert's answer
2022-05-23T23:18:42-0400

A Bernoulli differential equation


"\dfrac{dy}{dx}-\dfrac{1}{3x}y=\dfrac{x^2}{3}y^{-2}"

Substitute

"z=y^{1-(-2)}=y^3"

"\dfrac{dz}{dx}=3y^2\dfrac{dy}{dx}"

"\dfrac{dy}{dx}=\dfrac{1}{3y^2}\dfrac{dz}{dx}"

Then


"\dfrac{dz}{dx}-\dfrac{1}{x}z=x^2"

"IF=e^{\int (-1/x)dx}=\dfrac{1}{x}"

"\dfrac{1}{x}\dfrac{dz}{dx}-\dfrac{1}{x^2}z=x"


"\dfrac{d}{dx}(\dfrac{z}{x})=x"

"\dfrac{z}{x}=\dfrac{x^2}{2}+C"

"z=\dfrac{x^3}{2}+Cx"

"y^3=\dfrac{x^3}{2}+Cx"